OpenStudy (anonymous):

Solve the equation. Round the result to the nearest hundredth. Need help please !! 1) 21x + 60 = 72 2) 23 - 5x = 114x + 30 3) 44y - 18 = y + 17 4)12b - 93 =54b - 142

OpenStudy (anonymous):

on 1) 21x + 60 = 72 subtract 60 on both sides, gives 21x = 72-60 21x= 12 now divide both sides by 21, gives 21x/21= 12/21 x=12/21 = 4/7 or x=0.57 ...ans now try that to the next prob

OpenStudy (anonymous):

Thanks (:

OpenStudy (anonymous):

ok is everything clear now? do you get them better?

OpenStudy (anonymous):

we can try them here say prob 3) 44y - 18 = y + 17 can you put all the Y only in the left hand side?

OpenStudy (anonymous):

3) 44y - 18 = y + 17 subtract y on both sides,also add 18 onbothside,gives 44y - y = 17 +18 gives 43y = 35 now divide 43 on both sides, gives 43y/43 = 35/43 y= 35/43 use calculator y=0.81 ans...